In 2017, the Acme Company had 30 individuals with recordable injuries, four individuals with a recordable illness from an occupational exposure, but no fatalities. The company had 400 employees who worked 40-hour weeks with 2 weeks off during the year. Calculate the total injury and illness rate (IR) for the company in 2017 using the following formula: IR = N X 200,000 ÷ T (N = number of injuries illnesses, and fatalities; Y = total hours worked by all employees during the period in question). Discuss how the calculated IR might be used by OSHA in determining programmed inspection schedules and how the company might use the IR to improve safety at the facility.
Answer #1

Total injury and illness rate(IR) is given by ,

IR=Number of injuries,illnesses,fatalities *200000 /Number of hours worked by all the employees

=(30+4) *200000 /(400*40*50)

=34*200000/800000

=34/4

=8.5

Based on the industry of Acme and the number of employees ,OSHA has a predetermined value for the incident rate.Suppose ,the IR exceeds the cut off or the allowed limit ,then OSHA schedules inspection of the firms.For example ,for a firm in the manufacturing industry the maximum allowable IR rate is 7.5 .Any value above this will alert OSHA and they would start the inspection process.

From the perspective of the firm ,if their IR value exceeds the allowable limit ,they need to enhance their safety measures and ensure that the IR value is brought down.